OpenStudy (aaronandyson):

Prove that cosecA/cosecA - 1 + cosec A/cosecA + 1 = 2sec^2A @welshfella @loser66

OpenStudy (aaronandyson):

@TheSmartOne

OpenStudy (aaronandyson):

@perl

OpenStudy (aaronandyson):

@jim_thompson5910

OpenStudy (aaronandyson):

@zepdrix

TheSmartOne (thesmartone):

Ah, I'm not able to think properly right now and I'm tired. :/ Maybe @zepdrix @Nnesha @jim_thompson5910 @perl can help you better :)

OpenStudy (perl):

Is this copied correctly? $$ \Large \frac{\csc A}{\csc(A )- 1} + \frac{\csc A}{\csc(A )+ 1} = 2\sec^2A $$

OpenStudy (aaronandyson):

Yes, @perl @zepdrix and no @Nnesha

OpenStudy (aaronandyson):

Now you're correct @Nnesha

OpenStudy (perl):

You can start with either side, lets go with the left side since we can simplify that $$\large{ \frac{\csc A}{\csc(A )- 1} + \frac{\csc A}{\csc(A )+ 1} \\ ~\\= \color{red}{\frac{\csc (A)-1}{\csc(A )- 1}}\cdot \frac{\csc A}{\csc(A )- 1} + \frac{\csc A}{\csc(A )+ 1}\cdot \color{red}{\frac{\csc (A)+1}{\csc(A )+ 1}} }$$

OpenStudy (perl):

$$\large { \frac{\csc A}{\csc(A )- 1} + \frac{\csc A}{\csc(A )+ 1} \\ ~\\= \normalsize \left(\color{red}{\frac{\csc (A)-1}{\csc(A )- 1}}\right) \cdot \left(\frac{\csc A}{\csc(A )- 1}\right) + \left(\frac{\csc A}{\csc(A )+ 1}\right) \cdot \left(\color{red}{\frac{\csc (A)+1}{\csc(A )+ 1}} \right) }$$

OpenStudy (aaronandyson):

The thing is not clear I mean I can't see the latex Its like I see half only.

OpenStudy (perl):

can you see it now?

OpenStudy (aaronandyson):

Nope :(

OpenStudy (perl):

you can try refreshing

OpenStudy (aaronandyson):

I did but It ain't working :/

OpenStudy (perl):

youre saying your latex does not show?

OpenStudy (perl):

can you take a screenshot of it

OpenStudy (aaronandyson):

http://prntscr.com/75dbgk

OpenStudy (perl):

thats weird

OpenStudy (perl):

this is what i see http://prntscr.com/75dblz

OpenStudy (perl):

anyways, its not a problem :)

OpenStudy (aaronandyson):

Can you type it out please?

TheSmartOne (thesmartone):

Hmm, seems like you are using a smaller screen thus the issues.

OpenStudy (perl):

$${ \frac{\csc A}{\csc(A )- 1} + \frac{\csc A}{\csc(A )+ 1} \\ ~\\= \left(\color{red}{\frac{\csc (A)-1}{\csc(A )- 1}}\right) \cdot \left(\frac{\csc A}{\csc(A )- 1}\right) \\+ \left(\frac{\csc A}{\csc(A )+ 1}\right) \cdot \left(\color{red}{\frac{\csc (A)+1}{\csc(A )+ 1}} \right) }$$

OpenStudy (perl):

is that better?

OpenStudy (aaronandyson):

YES!

OpenStudy (perl):

ok great :)

OpenStudy (perl):

note that what i multiplied by was not arbitrary, we want a common denominator

OpenStudy (perl):

and its a valid multiplication, since it is equivalent to multiplying by 1. therefore i did not change the expression

OpenStudy (aaronandyson):

But I was told to do this cosec A/cosecA-1 x cosec A +1/cosec A + 1

OpenStudy (perl):

thats essentially the same thing, just skipping the step of common denominator

OpenStudy (perl):

$$ \Large{ \frac a b + \frac c d \\~\\ = \frac{a\cdot d + b\cdot c }{b\cdot d} }$$

OpenStudy (aaronandyson):

What next?

OpenStudy (perl):

$$ \Large{ \frac A B + \frac C D = \frac{AD + BC }{BD} }$$

OpenStudy (aaronandyson):

???

OpenStudy (perl):

do you agree with that rule

OpenStudy (aaronandyson):

Sure.

OpenStudy (perl):

$$ { \frac{\csc A}{\csc(A )- 1} + \frac{\csc A}{\csc(A )+ 1} \\ ~\\ = \frac{\csc A \cdot ( \csc(A )+ 1) + \csc A \cdot (\csc(A )- 1) }{(\csc(A )- 1)~(\csc(A )+ 1)} }$$

OpenStudy (aaronandyson):

???

OpenStudy (aaronandyson):

Is it like 2csc^2A/cot^2A?

OpenStudy (perl):

$$ { \frac{\csc A}{\csc A - 1} + \frac{\csc A}{\csc A + 1} \\ ~\\ = \frac{\csc A \cdot ( \csc A+ 1) + \csc A \cdot (\csc A - 1) }{(\csc A - 1)~(\csc A + 1)} \\~\\ = \frac{\csc^2 A + \csc A + \csc^2 A- \csc A }{(\csc A - 1)~(\csc A + 1)} \\~\\ = \frac{2\csc^2 A }{(\csc A - 1)~(\csc A + 1)} \\~\\ = \frac{2\csc^2 A }{\csc^2 A - \csc A + \csc A - 1} \\~\\ = \frac{2\csc^2 A }{\csc^2 A - 1} }$$

OpenStudy (aaronandyson):

csc^2A - 1 = cot^2A?

OpenStudy (perl):

right :)

OpenStudy (perl):

$$ { \frac{\csc A}{\csc A - 1} + \frac{\csc A}{\csc A + 1} \\ ~\\ = \frac{\csc A \cdot ( \csc A+ 1) + \csc A \cdot (\csc A - 1) }{(\csc A - 1)~(\csc A + 1)} \\~\\ = \frac{\csc^2 A + \csc A + \csc^2 A- \csc A }{(\csc A - 1)~(\csc A + 1)} \\~\\ = \frac{2\csc^2 A }{(\csc A - 1)~(\csc A + 1)} \\~\\ = \frac{2\csc^2 A }{\csc^2 A - \csc A + \csc A - 1} \\~\\ = \frac{2\csc^2 A }{\csc^2 A - 1} \\~\\ = \frac{2\csc^2 A }{\cot^2 A } \\~\\ =\Large 2\cdot \frac{\frac{1}{\sin ^2 A } }{\frac 1 {1\tan^2 A} } \\~\\ =\large 2\cdot \frac{1}{\sin^2 A } \cdot \frac{\tan^2 A }{1} \\~\\ =\large 2\cdot \frac{1}{\sin^2 A } \cdot \frac{\sin^2 A }{\cos^2 A } \\~\\ =\large 2\cdot \frac{1}{\cancel \sin^2 A } \cdot \frac{\cancel \sin^2 A }{\cos^2 A } \\~\\ =\large 2\cdot \frac{1}{\cos^2 A } \\~\\ =\large 2\cdot \sec^2 A }$$

OpenStudy (aaronandyson):

Thanks, can u help me more?